A piece of fuse wire melts when the current is 5 A. The energy produced then is 1 J/s. The resistance of the fuse in ohm is
A. 0.04 B. 0.1 C. 0.5 D. 10

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

P = i*i*r

1=25*r

r=1/25

Regards
Vikas TU
3 years ago
Dear student 
Option A will be crrect ans...................................
Good Luck …....................................